Problem 13, page 182:

3(2x+6)-5(x+8)=x-22

Found 2 solutions by EdenWolf, ikleyn:
Answer by EdenWolf(517)
Answer by ikleyn(52780)

.
3(2x+6)-5(x+8)=x-22,

6x + 18 - 5x - 40 = x -22,

x - 22 = x - 22.

This is an identity.  Any value of  x  satisfies this identity.

Hence,  any value of  x  satisfies the original equation.
It has infinitely many solution:  all real numbers.